A⃗ =(7.2)i^+(6.7)j^+(−3.3)k^

A→=(7.2)i^+(6.7)j^+(−3.3)k^

B⃗ =(3.7)i^+(−3.5)j^+(−8.7)k^

B→=(3.7)i^+(−3.5)j^+(−8.7)k^

Find the value of the x

x-component of 2A⃗ ×B⃗ 



1
Expert's answer
2021-09-10T17:21:14-0400

"x-component\ of \ (2\vec{A} \times \vec{B})=(13.4\cdot(-8.7)-(-6.6\cdot(-3.5)))\vec{i}=93.48\vec{i}"